OpenStudy (anonymous):

I need help finding out what equations are perpendicular and parallel. Can someone help me? 3x+2y=7 and 3x-2y=7 3x+2y=7 and 3x-1/2y=7 3x+2y=7 and -1/3x+2y=7 3x+2y=7 and -1/3x+1/2y=7 3x+1/2y=7 and 1/3x+2y=7 3x+2y=7 and 2x+4/3y=7

OpenStudy (smurfy14):

put in y=mx+b form for each one and if the two m's are the same, they are parallel and if they are opposite it is perpendicular example if first equation was y=3x+3 and second one was y=-1/3x+3 it would be perpendicular